obtain the semi-empirical mass formula for a nuclide of mass number A and atomic C number Z.
1
Expert's answer
2020-04-27T10:13:08-0400


The mass of an atomic nucleus, for N neutrons, Z protons and therefore A = N + Z nucleons, is given by

m = Zmp + N mn − EB ( N , Z ) / c2

where mp and mn are the rest mass of a proton and a neutron, respectively and EB is the binding energy of the nucleus. The semi-empirical mass formula states the binding energy is:

EB = aVA − aSA2 / 3 − aC * Z ( Z − 1 ) / A1 / 3 − aA * ( A − 2 Z) 2/ A − δ ( N , Z )


The δ ( N , Z ) term, depending on N and δ0 = aPAkp , is either zero or ± δ0, where δ0  =  aPAkp for some exponent kP .

Each of the terms in this formula has a theoretical basis. The coefficients aV, aS ,aC , aA are determined empirically; while they may be derived from experiment, they are typically derived from least squares fit to contemporary data.